Please answer these question:

_ An aqueous solution of a weak acid, HWk (pKa=5.25), has a pH of 4.53 . calculate the molarity of the weak acid.

_ A 0.15M solution of a weak base was found to have a pH =9.26 . calculate the Kb for this base.
Answer #1

Answer Let Go is molarity of weak ald ICE hable HA CON = Home Initial Change equilibrium ka = [HT] [A] MA) ka = x² Now pH - L5.62 x 10-6 = (2-96 6.- (2.98x10-5) ↑ 6-(2.95X10-5) = 1.5485 710-4 = 1.84 X10 M i nitial concentration of weak acid Ile table2. 21 X{o-9 | Ki = (A-o/Kz 8-1) - 51.0 - (-s-0IX28 -1) - = १ - 5-0) ४८४ | = = E-HQ] r-51.0 २८ = पत्र